OpenStudy (anonymous):

how to factor: 1. x^3+8x^2+16x=0 2. x^3+x^2-25x=25

zepdrix (zepdrix):

Ok the first one is pretty straight forward, if we factor an \(\large x\) out of each term, we'll be left with a quadratic in the brackets.\[\large x(x^2+8x+16)=0\]Understand that step?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

So from here, we want numbers that `multiply` to give us \(\large 16\), and those same numbers when we `add` them should give us \(\large 8\). Those are the two numbers that will give us our factors. \(\large 16\) and \(\large 1\). \(\large 16 \times 1=16\) But when we add them \(\large 16+1\neq8\). So those values won't work out. Let's try some other multiples of 16. How about \(\large 4\) and \(\large 4\), will those work?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\(\large 4\times4=16\) and \(\large 4+4=8\) Cool! So we can write our factors as two binomials with x and 4. \[\large x(x^2+8x+16)=0 \qquad \rightarrow \qquad x(x+4)(x+4)=0\] Good? :)

OpenStudy (anonymous):

yes so then i set both binomials equal to zero?

zepdrix (zepdrix):

If the problem is asking you to find the `roots` or `solutions` then yes, you would use the `Zero Factor Property` ~ Set each factor equal to 0, and solve for x separately.\[\large x=0 \qquad \qquad x+4=0 \qquad \qquad x+4=0\] But if the directions just say "factor this", then we were done with our last step.

OpenStudy (anonymous):

ok thx for the help

zepdrix (zepdrix):

np c:

OpenStudy (anonymous):

ya math is my worst subject thats why i always need help with my math homework

OpenStudy (anonymous):

what about the 2nd problem?

zepdrix (zepdrix):

Ah that stinks c: we all have different strengths! Hmm this next problem I'm a little confused on... I always forget how to factor cube functions... thinkinggg c:

OpenStudy (anonymous):

ya it does since i want to go into medical school and have to major in it and its ok

zepdrix (zepdrix):

Oh for this next one I think we can solve by grouping. Lemme try real quick before I explain it.

OpenStudy (anonymous):

ok sure

zepdrix (zepdrix):

\[\large x^3+x^2-25x=25\]Ok let's start by subtracting 25 from each side,\[\large \color{royalblue}{x^3+x^2}-25x-25=0\]From here, we'll factor out an x^2 from each of the blue terms. \[\large x^2(\color{royalblue}{x+1})-25x-25=0\]

zepdrix (zepdrix):

\[\large x^2(x+1)\color{orangered}{-25x-25}=0\]Then let's factor a -25 out of each orange term.\[\large x^2(x+1)-25(\color{orangered}{x+1})=0\]

zepdrix (zepdrix):

Understand those steps? Try to digest that a moment c: Lemme know if you're confused so far.

OpenStudy (anonymous):

i understand so far

zepdrix (zepdrix):

This next step is a littleeeee tricky. We have 2 terms. They share a factor. Though, it's not a number or simply x^2 that we're factoring out this time. We want to factor out \(\large (x+1)\) from each term.\[\large \color{#F35633}{x^2}(x+1)\color{#F35633}{-25}(x+1)=0 \qquad \rightarrow \qquad (x+1)\left(\color{#F35633}{x^2-25}\right)=0\] Hopefully the color coding helps you see where everything went.

zepdrix (zepdrix):

We have one more simplification we can do. I'll wait and see if you understand that part first though.

OpenStudy (anonymous):

i understand it

zepdrix (zepdrix):

For the last step, we'll need to recall this identity. \(\large a^2-b^2 \qquad = \qquad (a-b)(a+b)\) It tells us how to factor the `difference of squares`. In our problem we have,\[\large (x+1)\left(x^2-25\right)=0\]Which we can write as,\[\large (x+1)\left(x^2-5^2\right)=0\] Because we recognize that \(\large 25=5^2\). Now we have the difference of squares in our right set of brackets. Can you see how it will factor?

OpenStudy (anonymous):

(x-5)(x+5)?

zepdrix (zepdrix):

Yes very good :)

zepdrix (zepdrix):

Which gives us a final answer of, \[\large (x+1)(x+5)(x-5)=0\] Yay team!

OpenStudy (anonymous):

yay thanks for the help @zepdrix